OpenStudy (anonymous):

Write an algebraic equation for the following: Martha’s age, M, is the same as Bob’s age, B, decreased by 9.

OpenStudy (anonymous):

M - B = 9 M = 9 - B M = B / 9 M = B - 9 answers

OpenStudy (anonymous):

M = B - 9